# Number of local minima in discrete-time fractional Brownian motion

### v2

@@ Abstract
where $\Phi=(\phi_{0},\dots,\phi_{N-1})$}, and ${\bf R}=(\rho_{ij})$} is the correlation matrix, whose elements depend only on the distnance between increments, $\rho_{ij}=\rho_{0|i-j|}$ and
@@
is linear for large $N$ as in Markovian case [21]. Meanwhile its prefactor is modified, but it depends only on the nearest-neighbor increment correlation $\rho_{01}$. In fact, it is the same as for the so-called autoregressive model AR(1) [54, 55], which is a discretized Ornstein-Uhlenbeck process [56], and whose correlation matrix is given by:
@@
at leading order in $N$ (see Supplemental Material (SM)  [^2] See Supplemental Material at [url] for detailed calculations and numerical checks of asymptotics for detailed calculations and numerical checks of asymptotics, as well as for definitions of coefficients $c_{H}$ and $b_{3/4}$). We note that the asymptotic scaling exponents of Eq. (6) also follow from Ref. [58], which considers the variance of the number of zero crossings of fGn in a unit interval. Because the fGn $\phi_{i}$ is a discrete derivative of the dicrete-time fBm $X_{i}$, $\phi_{i}=X_{i+1}-X_{i}$, the number of local minima $m_{N}$ of fBm is equal to the number of zero crossings of fGn in a unit interval multiplied by $N/2$ and their variances are related by the factor $N^{2}/4$, respectively. However, in the regime $H>3/4$, which is a key analytic result, Ref. [58] contains an error in the prefactor; our result (6) corrects this prefactor (see SM for two independent proofs). Importantly, this behavior comes from the long-range increment correlations $\rho_{0k}$. Given that $\rho_{0k}\propto k^{2H-2}$ for $k\rightarrow\infty$ and that the correlation function $\rho_{0k}$ is squared in $\text{Var}(m_{N})\propto\sum_{k=2}^{N-2}(N-k-1)\rho_{0k}^{2}$, the correlations become non-summable when 2(2H2)=12(2H-2)=1, i.e. at $H=3/4$, which is not the usual threshold value $H=1/2$ that separates fBm sub- and superdiffusion  [^3] On the level of a path-integral representation, the fBm action involves a kernel with decay exponent 22H2-2H [80]; here the correlations of the minima involve the squared correlations of fBm increments, which are expected to be controlled by an effective kernel with exponent 44H4-4H (which becomes marginal, equal to $1$, at $H=3/4$)..
@@
The strong correlations of $\{\phi_{i}\}$ leading to the anomalously large (superlinear) fluctuations of $\text{Var}(m_{N})$ have a fundamental consequence for the distribution of $m_{N}$, for which the CLT breaks down. We show below this breakdown analytically and confirm it numerically through simulations (see Fig. 2(a))  [^4] The breakdown was also found in the continuous setting of the counterpart zero-crossing problem for fBm studied by Slud, who showed that the CLT fails for $H>3/4$ by representing the number of zero crossings as a second-order Wiener–Itô integral [72, 73].. In the analytic approach, we use the methods Ref. [61] for functionals of Gaussian correlated variables, which allow us to go beyond one-time distribution and to show the convergence of $m_{N}$ to the processes belonging to two different universality classes (Rosenblatt and Brownian motion), separated through the threshold value $H=3/4$.
@@
Because the rank $2$ terms drive the CLT breaking, we are going now to analyse Eq. (9), by spanning the quadratic part $M_{i}^{\circ}\equiv M_{i}-\langle M_{i}\rangle$ on the orthogonal vectors of the complete quadratic Wick basis
@@
$Z_{K}(t)=\frac{m_{[Kt]}-\langle m_{[Kt]}\rangle}{\sqrt{m_{K}}},$ (21)

### v4

@@ Abstract
where $\Phi=(\phi_{0},\dots,\phi_{N-1})$}, and ${\bf R}=(\rho_{ij})$} is the correlation matrix, whose elements depend only on the distance between increments, $\rho_{ij}=\rho_{0|i-j|}$ and
@@
is linear for large $N$ as in Markovian case [21]. Meanwhile its prefactor is modified, but it depends only on the nearest-neighbor increment correlation $\rho_{01}$. In fact, it is the same as for the so-called autoregressive model AR(1) [60, 61], which is a discretized Ornstein-Uhlenbeck process [62], and for the model with the correlation matrix:
@@
at leading order in $N$ (see Supplemental Material (SM)  [^2] See Supplemental Material at [url] for detailed calculations and numerical checks of asymptotics, which includes Refs. [59, 64, 38, 21, 22, 60, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75] for detailed calculations and numerical checks of asymptotics, as well as for definitions of coefficients $c_{H}$ and $b_{3/4}$). We note that the asymptotic scaling exponents of Eq. (6) also follow from Ref. [65], which considers the variance of the number of zero crossings of fGn in a unit interval. Because the fGn $\phi_{i}$ is a discrete derivative of the dicrete-time fBm $X_{i}$, $\phi_{i}=X_{i+1}-X_{i}$, the number of local minima $m_{N}$ of fBm is equal to the number of zero crossings of fGn in a unit interval multiplied by $N/2$ and their variances are related by the factor $N^{2}/4$, respectively. However, in the regime $H>3/4$, which is a key analytic result, Ref. [65] contains an error in the prefactor; our result (6) corrects this prefactor (see SM for two independent proofs). Importantly, this behavior comes from the long-range increment correlations $\rho_{0k}$. Given that $\rho_{0k}\propto k^{2H-2}$ for $k\rightarrow\infty$ and that the correlation function $\rho_{0k}$ is squared in $\text{Var}(m_{N})\propto\sum_{k=2}^{N-2}(N-k-1)\rho_{0k}^{2}$, the correlations become non-summable when 2(2H2)=12(2H-2)=-1, i.e. at $H=3/4$, which is not the usual threshold value $H=1/2$ that separates fBm sub- and superdiffusion  [^3] On the level of a path-integral representation, the fBm action involves a kernel with decay exponent 22H2-2H [77]; here the correlations of the minima involve the squared correlations of fBm increments, which are expected to be controlled by an effective kernel with exponent 44H4-4H (which becomes marginal, equal to $1$, at $H=3/4$)..
@@
The strong correlations of $\{\phi_{i}\}$ leading to the anomalously large (superlinear) fluctuations of $\text{Var}(m_{N})$ have a fundamental consequence for the distribution of $m_{N}$, for which the central limit theorem (CLT) breaks down. We show below this breakdown analytically and confirm it numerically through simulations (see Fig. 2(a)). In the analytic approach, we use the methods Ref. [66] for functionals of Gaussian correlated variables, which allow us to go beyond one-time distribution and to show the convergence of $m_{N}$ to the processes belonging to two different universality classes (Rosenblatt and Brownian motion), separated through the threshold value $H=3/4$. Note that, in the continuous setting, as outlined in the review by Kratz (Th.2.27) [78], the limiting process of the related problem of zero crossings’ number is given in the supercorrelated regime through a form involving implicitly the correlation function and is expressed through the multiple Wiener integral of the second order. This fundamental mathematical object, originally found by Slud [69], represents the Rosenblatt process, but practically is of little use. Here, in the discrete setting, we construct the Rosenblatt process in closed form with explicit analytic parametrisation.
@@
[…]
Because the rank $2$ terms drive the CLT breaking, we are going now to analyse Eq. (9), by spanning the quadratic part $M_{i}^{\circ}\equiv M_{i}-\langle M_{i}\rangle$ on the linearly independent vectors of the complete quadratic Wick basis
@@
$Z_{K}(t)=\frac{m_{[Kt]}-\langle m_{[Kt]}\rangle}{\sqrt{\mathrm{Var}(m_{K})}},$ (21)
@@
Data availability. The data that support the findings of this article are openly available [83].